# The Monge optimal transport barycenter problem

### v1

@@ 6.1 Choices for the functional spaces $\mathcal{F}$ and $\mathcal{G}$
The methodology is not fully specified until we select the functional spaces $\mathcal{F}$ and $\mathcal{G}$ to use. This section describes some practical choices and their consequences, including the choices that we have made for the simulations in Section 9. A more complete exploration of data-adapted functional spaces is beyond the scope of this article, it will be pursued elsewhere.
Even though the notion of independence is symmetric in $y$ and $z$, the consequences of restricting the two functional spaces $\mathcal{G}$ and $\mathcal{F}$ enforcing independence through (4) are conceptually quite different. While the richness of $\mathcal{F}$ relates to the level of resolution of the dependence of $x$ on $z$, the richness of $\mathcal{G}$ specifies the level of detail with which the distributions $\rho(x|z)$ are captured for each value of $z$. Imagine temporarily that $\mathcal{F}$ is rich enough to capture all functions $f(z)$ that may be required, and concentrate on the effect of making different choices for $\mathcal{G}$. Since for any $g\in\mathcal{G}$, the condition
[…]
must hold for any $f\in\mathcal{F}$ with zero mean –and $\mathcal{F}$ is by hypothesis rich enough– it follows that
$\int g(y)\ d\rho(y|z)\perp\!\!\!\perp z,$
i.e. the conditional expectation $\bar{g}(z)$ is in fact a constant $\bar{g}$ independent of $z$.
When $X=R^{d}$, the simplest choice for $\mathcal{G}$ is the space of linear functions, spanning the columns of a matrix $G$ with the $d$ independent functions $y^{l},\ l\in[1,\ldots,d]$. It follows from the argument above that the conditional mean of $y$ is independent of $z$. On the other hand, from (12), $X(y,z)=y+h(z)$, where $h(z)=2\lambda\ \sum_{k=1}^{K}\sigma_{k}\ f_{k}(z)\ v_{k}$ and $v_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\nabla_{y}g_{k}(y)$ is a constant, since $g_{k}(y)$ is linear. It follows from taking the conditional expectation of both sides that
$h(z)=\bar{x}(z)-\bar{y},$
so the procedure captures –and removes from $x$– the conditional mean $\bar{x}(z)$, i.e. it performs [nonlinear] regression, as in the “poorest man solution” of [1]. When the columns of $G$ span a general quadratic function of $y$, not only the conditional mean but also the conditional covariance matrix of $y$ is independent of $z$, and (12) implies that the relation between $x$ and $y=T(x,z)$ is linear (with $z$-dependent coefficients), as in the “poor man solution” of [1]. We typically start our experiments with a first run where the columns of $G$ span all quadratic functions of $y$, to capture the conditional mean and covariance matrix of $x$, leaving a more detailed characterization of $\rho(x|z)$ to subsequent runs.
One can of course extend the choices above and fill all columns of $G$ with externally provided functions, such as Hermite polynomials of a given degree. Similarly, we can use as columns of $F$ polynomials when $z\in R^{d}$ , trigonometric functions when $z$ is a periodic variable, such as the time of the day, and indicator functions when $z$ can only adopt a finite set of categorical values. Yet it is generally preferable to use a less parametric approach and let the data dictate the form of the functions to use. For our experiments, we have used a simple class of data-adapted spaces described in the appendix, where the columns of $F$ and $G$ are given by asymmetric kernel-like functions with column dependent center and bandwidths, a flexible and economic variation of reproducing kernel Hilbert spaces:
@@
Even though the $\{y_{i}\}$ evolve, the centers $\{y^{c}_{j}\}$ are fixed throughout stages of the procedure, so as to have a fixed functional space $\mathcal{G}$. Their cardinality does not need to match that of the $\{y_{i}\}$, in practice it is typically much smaller.
@@ 6.2 Choice of the penalization parameter
The penalization parameter $\lambda$ establishes a balance at the final $y=y_{*}$ between the gradients of the transportation cost function and the penalization term. It follows from (11) that if one would like the $\sigma_{k}\left(y_{*}\right)$ to have values smaller than $\sigma_{*}\ll 1$, one must adopt a parameter $\lambda$ satisfying
@@
where $\tilde{\sigma}_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\min(\sigma_{k},\sigma_{*})$ and $\overline{\|s\|^{2}}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n}\sum_{i=1}^{n}\|s_{i}\|^{2}$ for any $s=\{s_{i}\}_{i=1,...,n}$, $s_{i}\in R^{d}$. Based on this characterization of $\lambda$ at convergence, we adopt a state-dependent penalization parameter that evolves over algorithmic time:
@@
The addition of a small fraction of the variance to the numerator addresses the fact that $\nabla_{y}c=0$ at the onset of the algorithm, when $x=y$ (The variance is a natural reference value, since $\overline{\left\|\nabla_{y}c\left(x,y\right)\right\|^{2}}=\frac{1}{n}\sum_{i}\left\|x_{i}-y_{i}\right\|^{2}\leq\frac{1}{n}\sum_{i}\left\|x_{i}-\bar{x}\right\|^{2}=\hbox{var}(x)$.)
6.3 Learning rate
We minimize the objective function $L$ in (11) through gradient descent, with steps of the form
$y_{i}^{n+1}=y_{i}^{n}-\eta^{n}\left.\nabla_{y_{i}}L\right|_{y^{n}},$
determining the learning rates $\eta^{n}$ through back-tracking limited to a small interval:
  1. 1.
    At time $t_{n}$, pick an initial candidate $\eta$ through $\eta=\theta\ \eta^{n-1}$, where $\theta$ is only slightly larger than one, so as to explore an interval for $\eta^{n}$ that is only marginally larger than the accepted learning rate from the previous step.
  2. 2.
    Back-track from this $\eta$ through the Armijo-Goldstein algorithm: defining $G=\left.\nabla_{y}L\right|_{y^{n}}$, set $\eta^{n}=\eta$ if $L\left(y^{n}-\eta G\right)\leq L\left(y^{n}\right)-\kappa\eta\|G\|^{2},\hbox{ with }\ 0<\kappa<1$. Otherwise, set $\eta\rightarrow\tau\eta$, $10<\tau<$ 1 and repeat this step.
We have adopted for our numerical examples $\theta=1.1$ and $\kappa=\tau=\frac{1}{2}$.
@@ 6.4 Termination criterion
    In order to assign a value to $\sigma_{*}$, notice that $\sigma_{k}$ represents the empirical correlation between $f_{k}(z)$ and $g_{k}(y)$, which should be uncorrelated. It follows that a reference value for $\sigma_{*}$ is the standard deviation of the empirical correlation between two independent variables, which equals $\frac{1}{\sqrt{n}}$. We have adopted in our experiments
@@
Another, internal termination criterion starts a new stage once the current $y$ differ significantly from their values $y^{0}$ at the outset of the current stage, i.e. when
$\sum_{i=1}^{n}\left\|y_{i}-y_{i}^{0}\right\|^{2}>\delta\sum_{i=1}^{n}\left\|y^{0}_{i}-\bar{y^{0}}\right\|^{2},\quad 0<\delta<1.$
In our experiments, we have adopted $\alpha=0.0025$ and $\delta=0.1$,
@@ 6.5 Successive barycenter problems
The barycenter problem removes from $x$ any $z$-dependence detectable through the functional spaces $\mathcal{F}$ and $\mathcal{G}$, so the resulting $y=T(x,z)$ can still depend on $z$ in ways that $\mathcal{F}$ and $\mathcal{G}$ do not capture. For instance, if $\mathcal{G}$ consists only of quadratic function of $y$, just the conditional mean and covariance of $x$ are removed, leaving in $y$ any other $z$-dependent property of $\rho(x|z)$, such as higher moments or the distribution’s modality. Similarly, if $\mathcal{F}$ includes only functions of a subset of the $\{z^{l}\}$, $y$ may still depend on the remaining ones, if $\mathcal{F}$ includes only functions of the individual $\{z^{l}\}$, any non-additive dependence of $x$ on the $\{z_{l}\}$ will remain in $y$, and if the bandwidths of the functions in $\mathcal{F}$ are large, only long-scale trends are removed, leaving small-scale signals behind.
This suggests proceeding in $N_{s}$ stages: calling $y^{0}=x$, we compute in stage $l$ the barycenter of the $y^{l-1}$ over $z$, as captured through the functional spaces $\mathcal{F}^{l}$ and $\mathcal{G}^{l}$. The final $y^{N_{s}}$ are not samples of the barycenter of the original $x$, since the composition of optimal maps is not necessarily optimal. However, since we know how to invert each of the maps, we can still simulate and estimate by composition $\rho(x|z^{*})$ for any target $z^{*}$. This procedure resembles boosting [34], in which multiple models are trained sequentially so that each new problem removes further variability from the barycenter of the prior one.
@@ 6.6 Complexity
One major advantage of the new methodology is its efficiency, which makes it applicable to large data sets. This efficiently derives from formulating the independence conditions between $y$ and $z$ in terms of the uncorrelation between test functions and subsequently relaxing it to the vanishing of the first singular value of a matrix $A(y)$ whose rank does not depend on the sample size. Previous methods used kernel density estimation to formulate independence [5], requiring kernels where every data point acted as a center, yielding at least O$(n^{2})$ time complexity. By contrast, the new algorithm’s time complexity scales bilinearly with the number of samples ($n$) and the dimension of the data ($d_{x}$). Because of this, most of the numerical examples in Section 9 had running times on a laptop raging between a fraction of a second and a few seconds.
The algorithm’s operations can be broken into three different categories: those which are performed only once per run, those performed once per stage, and those performed at each descent step. Even though it is only the third category that determines the time complexity of the algorithm in practice, we analyze all three parts for completeness.
  1. 1.
@@
    Operations performed only once. Since the factors $z$ do not evolve through a run, the orthogonal matrix $Q_{z}$ is computed only once, making the algorithm’s complexity insensitive to the dimensionality of $z$. When using kernels, calculating $Q_{z}$ requires k-means clustering to determine the centers for $z$, which with a fixed maximum number of iterations requires O($n*d_{z}*m_{z}$) operations. Evaluating the kernel function also requires O($n*d_{z}*m_{z}$) operations. A standard singular value decomposition of the matrix $F\in\mathbb{R}^{n\times n_{z}}$ requires O($n*n_{z}^{2}$) steps where $n_{z}$ is a user’s provided input. Therefore, the number of operations performed only once scales as O($n*d_{z}*m_{z}$). For very large data-sets with high-dimensional factors $z$, this number can be further reduced by adopting state-of-the-art methodologies for finding the first few principal components of large matrices [35, 36], but we found no need for this in our experiments so far. Even for time-like factors $z$, which as discussed in the appendix require a number of centers that grows linearly with the extent of the time series analyzed, the corresponding matrix $F(z)$ is sparse, which keeps the complexity of its principal component analysis at $O(n)$, since the complexity scales not with the number of columns of $F$ (i.e. the number of centers) but with the number of non-zero elements in each row, which can be kept fixed using kernel functions with compact support.
  2. 2.
@@
    Operations performed once per stage. To recall, a new stage is started when the average squared distance between the current values of the $\{y_{i}\}$ and their values $\{y_{i}^{0}\}$ at the start of the current stage is larger than a prescribed fraction of the variance of the latter. At the beginning of each stage, the orthogonal matrix $Q_{y}$ needs to be computed. The same scaling arguments apply here as for the calculations of $Q_{z}$, yielding a total of O($n*d_{y}*m_{y}$) operations per stage.
  3. 3.
@@
    Operations performed at each descent step. The main loop iteration requires calculating derivatives of the cost function and of the penalty function. The complexity of the former is O($n*d_{y}*m_{y}$)). The latter requires calculating the gradient of the matrix $G(y)$ which, when using kernels, involves calculating the kernel in $y$ and its derivatives, with complexity O($n*d_{y}*m_{y}$) if the number of kernel centers is fixed. This is followed by a matrix multiplication which is O($n$). So overall each iteration performs O($n*d_{y}*m_{y}$) operations. Moreover, these operations decouple among the sample points, making them trivially parallelizable.
@@
One thing not captured by the complexity analysis above is the number of iterations required for convergence. Additionally, in practice one may adopt larger values of $n_{y}$ and $n_{z}$ for problems with more complex dependence between $x$ and $z$. Yet, for a fixed problem, the number of iterations should not depend on the number of data points $n$, an observation confirmed in our numerical experiments. We verify the algorithm’s complexity by plotting the time of the pre-calculations, stage calculations, average descent iteration and total time as $n$, $d_{x}$ and $d_{z}$ vary. In each case, each factor $z^{l}$ is a normal random variable with mean 00 and variance $0.25$ and $x$ is drawn from the $z$-dependent isotropic gaussian
@@
Figure 1 displays the data and barycenter for $d_{x}=d_{z}=1$ and figure 2 shows the running times for various values of $n$, $d_{x}$, and $d_{z}$. Each data point displayed is the median across 10 trials of the mean time spent in each portion of the algorithm. When the dependence on $n$ is being considered, $d_{x}$ and $d_{z}$ are both kept at 1. When either $d_{x}$ or $d_{z}$ are being varied, the other is kept at 1 and $n$ is kept at 2000. The first two rows use kernel based functions for both $F$ and $G$ while the third row uses only linear and quadratic functions of $y$ for $G$, which has the same complexity but with lower constants and thus is orders of magnitude faster. These experiments confirm that the complexity of all stages grow linearly with the number of samples and that the descent steps contribute most heavily to the total run time. Similarly we see that, as predicted, the dimension of $x$ increases linearly the time complexity of the main descent steps and the calculations per stage, and the dimension of $z$ increases linearly the complexity of the pre-calculations.
@@ 7 The optimal transport problem
The methodology developed for the optimal transport barycenter problem can be easily adapted to solve regular [Monge] optimal transport problems. In addition to the importance of the optimal transport problem per se, we will use it for conditional density estimation, as detailed in the next section. Optimal transport problems are simpler than the OTBP, as they involve only two distributions, a source $\rho_{0}$ and a target $\rho_{1}$:
@@
For the purpose of relating them to the barycenter problem, we introduce a binary covariate $z\in\{0,1\}$, and think of the source and target distributions as instances of the single conditional distribution $\rho(x|z)$:
$\rho_{0}(x)=\rho(x|z=0),\quad\rho_{1}(x)=\rho(x|z=1).$
Then the map $y=T(x,z)$ that solves the barycenter problem for $\rho(x|z)$, automatically provides the solution to (17) through
$Q(x)=T^{-1}\left(T(x,0),1\right),$
a standard result in interpolation displacement [3].
@@
The fact that there is only one, binary covariate $z$ simplifies our methodology considerably, since except for an arbitrary sign, there is only one function $f(z)$ with zero mean and norm one:
$f(z)\propto\begin{cases}\ \frac{1}{n_{0}}&\hbox{for }z=0\cr-\frac{1}{n_{1}}&\hbox{for }z=1\end{cases}.$
Then the barycenter problem reduces to
@@
Other than the simplifying facts that we do not need to update $a$ and $b$ and that the matrix $Q_{z}$ consists of a single column, the procedure to solve (18) follows the same steps as the one for the full barycenter problem (11).
In order to bypass the barycenter $\mu$ in the procedure above, finding a map $Q(x)=T(x,0)$ that pushes forward $\rho_{0}$ to $\rho_{1}$ directly, use the same objective function $L$ in (18), but minimize it only over the $y_{i}$ with corresponding $z_{i}=0$, i.e. over $T(x,0)$, leaving the remaining $y_{i}$ fixed at $x$, i.e. setting $T(x,1)=x$. This enforces the condition that $T(x,0)\#\rho_{0}=\rho_{1}$, since the final
$y_{i}^{*}=\begin{cases}T\left(x_{i},0\right)&\hbox{for }z_{i}=0\cr x_{i}&\hbox{for }z_{i}=1\end{cases}$
must be independent of $z$. This procedure, while lacking the symmetry of the prior one with respect to $\rho_{0}$ and $\rho_{1}$, is far more straightforward. In particular, it is very well-suited for density estimation.

### v2

@@ 6 Implementation
We minimize the objective function $L$ in (10) through gradient descent: $y_{i}^{n+1}=y_{i}^{n}-\eta^{n}\left.\nabla_{y_{i}}L\right|_{y^{n}}$, back-tracking the learning rate $\eta^{n}$ from $\theta\eta^{n+1}$, $\theta>1$, through the Armijo-Goldstein rule [1, 2].
@@ 6.1 Choices for the functional spaces $\mathcal{F}$ and $\mathcal{G}$
The methodology is not fully specified until we select the functional spaces $\mathcal{F}$ and $\mathcal{G}$. When $X=R^{d}$, the simplest choice for $\mathcal{G}$ is the space of linear functions, enforcing independence between any $f(z)$, $f\in\mathcal{F}$ and the conditional mean of $y$. Then, from (11), $X(y,z)=y+h(z)$, where $h(z)=2\lambda\ \sum_{k=1}^{K}\sigma_{k}\ f_{k}(z)\ v_{k}$ and $v_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\nabla_{y}g_{k}(y)$ is a constant, since $g_{k}(y)$ is linear. It follows that $h(z)=\bar{x}(z)-\bar{y}$, so for a rich enough space $\mathcal{F}$, the procedure captures –and removes from xx– the conditional mean $\bar{x}(z)$, i.e. it performs [nonlinear] regression, as in the “poorest man solution” of [3]. When the columns of $G$ span a general quadratic function of $y$, not only the conditional mean but also the conditional covariance matrix of $y$ is independent of $z$, and (11) implies that the relation between $x$ and $y=T(x,z)$ is linear (with $z$-dependent coefficients), as in the “poor man solution” of [3]. We typically start our experiments with a first run where the columns of $G$ span all quadratic functions of $y$, to capture the conditional mean and covariance matrix of $x$, leaving a more detailed characterization of $\rho(x|z)$ to subsequent runs.
One could extend the choices above and fill all columns of $G$ with externally provided functions, such as Hermite polynomials of a given degree. Similarly, we could use as columns of $F$ polynomials when $z\in R^{d}$ , trigonometric functions when $z$ is a periodic variable and indicator functions when $z$ can only adopt a finite set of categorical values. Yet it is generally preferable to use a less parametric approach and let the data dictate the form of the functions to use. For our experiments, we have used a simple class of data-adapted spaces described in the appendix, where the columns of $F$ and $G$ are given by asymmetric kernel-like functions with column dependent center and bandwidths, a flexible and economic variation of reproducing kernel Hilbert spaces:
@@
Even though the $\{y_{i}\}$ evolve, the centers $\{y^{c}_{j}\}$ are fixed throughout stages of the procedure, so as to have a fixed functional space $\mathcal{G}$. Their cardinality does not need to match that of the $\{y_{i}\}$, it is typically much smaller.
@@ 6.2 Choice of the penalization parameter
The penalization parameter $\lambda$ establishes a balance at the final $y=y^{*}$ between the gradients of the transportation cost function and the penalization term. It follows from (10) that, in order for the $\sigma_{k}\left(y^{*}\right)$ to have values smaller than $\sigma_{*}\ll 1$, the parameter $\lambda$ must satisfy
@@
where $\tilde{\sigma}_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\min(\sigma_{k},\sigma_{*})$ and $\overline{\|s\|^{2}}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{1}{n}\sum_{i=1}^{n}\|s_{i}\|^{2}$. Based on this characterization of $\lambda$ at convergence, we adopt a state-dependent penalization parameter that evolves over algorithmic time:
@@
The addition of a small fraction of the variance to the numerator addresses the fact that $\nabla_{y}c=0$ at the onset of the algorithm, when $x=y$. The variance is a natural reference value, since $\overline{\left\|\nabla_{y}c\left(x,y\right)\right\|^{2}}=\frac{1}{n}\sum_{i}\left\|x_{i}-y_{i}\right\|^{2}\leq\frac{1}{n}\sum_{i}\left\|x_{i}-\bar{x}\right\|^{2}=\hbox{var}(x)$.
@@ 6.3 Termination criterion
    In order to assign a value to $\sigma_{*}$, notice that $\sigma_{k}$ represents the empirical correlation between $f_{k}(z)$ and $g_{k}(y^{*})$, which should be uncorrelated for all $f\in\mathcal{F}$, $g\in\mathcal{G}$. It follows that a reference value for $\sigma_{*}$ is the standard deviation of the empirical correlation between two independent variables, which equals $\frac{1}{\sqrt{n}}$. We have adopted in our experiments
@@
Another, internal termination criterion starts a new stage once the current $y$ differ significantly from their values $y^{0}$ at the outset of the current stage, i.e. when $\sum_{i=1}^{n}\left\|y_{i}-y_{i}^{0}\right\|^{2}>\delta\sum_{i=1}^{n}\left\|y^{0}_{i}-\bar{y^{0}}\right\|^{2},\quad 0<\delta<1$. We have adopted in our experiments $\alpha=0.0025$ and $\delta=0.1$.
@@ 6.4 Successive barycenter problems
Since the barycenter problem removes from $x$ any $z$-dependence detectable through the functional spaces $\mathcal{F}$ and $\mathcal{G}$, $y=T(x,z)$ can still depend on $z$ in ways that $\mathcal{F}$ and $\mathcal{G}$ do not capture. For instance, if $\mathcal{G}$ consists only of quadratic function of $y$, just the conditional mean and covariance of $x$ are removed, leaving in $y$ any other $z$-dependent property of $\rho(x|z)$, such as higher moments or the distribution’s modality. Similarly, if $\mathcal{F}$ includes only functions of a subset of the $\{z^{l}\}$, $y$ may still depend on the remaining ones, if $\mathcal{F}$ includes only functions of the individual $\{z^{l}\}$, any non-additive dependence of $x$ on the $\{z_{l}\}$ will remain in $y$, and if the bandwidths of the functions in $\mathcal{F}$ are large, only long-scale trends are removed, leaving small-scale signals unresolved.
This suggests proceeding in $N_{s}$ stages: calling $y^{0}=x$, we compute in stage $l$ the barycenter of the $y^{l-1}$ over $z$, as captured through the functional spaces $\mathcal{F}^{l}$ and $\mathcal{G}^{l}$. The final $y^{N_{s}}$ are not samples of the barycenter of the original $x$, since the composition of optimal maps is not necessarily optimal. However, since we know how to invert each of the maps, we can still simulate and estimate by composition $\rho(x|z_{*})$ for any target $z_{*}$. This procedure resembles boosting [35], in which multiple models are trained sequentially so that each new problem removes further variability from the barycenter of the prior one.
@@ 6.5 Complexity
[…]
@@
Since the factors $z$ do not evolve through a run, the orthogonal matrix $Q_{z}$ is computed only once, making the algorithm’s complexity insensitive to the dimensionality of $z$. When using kernels, calculating $Q_{z}$ requires k-means clustering to determine the centers for $z$, which with a fixed maximum number of iterations requires O($n*d_{z}*m_{z}$) operations. Evaluating the kernel function also requires O($n*d_{z}*m_{z}$) operations. A standard singular value decomposition of the matrix $F\in\mathbb{R}^{n\times n_{z}}$ requires O($n*n_{z}^{2}$) steps where $n_{z}$ is a user’s provided input. Therefore, the number of operations performed only once scales as O($n*d_{z}*m_{z}$). For very large data-sets with high-dimensional factors $z$, this number can be further reduced by adopting state-of-the-art methodologies for finding the first few principal components of large matrices [36, 37].
@@
To recall, a new stage is started when the average squared distance between the current values of the $\{y_{i}\}$ and their values $\{y_{i}^{0}\}$ at the start of the current stage is larger than a prescribed fraction of the variance of the latter. At the beginning of each stage, the orthogonal matrix $Q_{y}$ needs to be computed. The same scaling arguments apply here as for the calculations of $Q_{z}$, yielding a total of O($n*d_{y}*m_{y}$) operations per stage.
@@
The main loop iteration requires calculating derivatives of the cost function and of the penalty function. The complexity of the former is O($n*d_{y}*m_{y}$)). The latter requires calculating the gradient of the matrix $G(y)$ which, when using kernels, involves calculating the kernel in $y$ and its derivatives, with complexity O($n*d_{y}*m_{y}$) if the number of kernel centers is fixed. This is followed by a matrix multiplication which is O($n$). So overall each iteration performs O($n*d_{y}*m_{y}$) operations, which decouple among the sample points, making them trivially parallelizable.
@@
Not captured by the complexity analysis above is the number of iterations required for convergence. Additionally, in practice one may adopt larger values of $n_{y}$ and $n_{z}$ for problems with more complex dependence between $x$ and $z$. Yet, for a fixed problem, the number of iterations should not depend on the number of data points $n$, an observation confirmed in our numerical experiments. We verify the algorithm’s complexity by plotting the time of the pre-calculations, stage calculations, average descent iteration and total time as $n$, $d_{x}$ and $d_{z}$ vary. In each case, each factor $z^{l}$ is a normal random variable with mean 0 and variance $0.25$ and $x$ is drawn from the $z$-dependent isotropic gaussian
@@
Figure 1 displays the data and barycenter for $d_{x}=d_{z}=1$ and figure 2 shows the running times for various values of $n$, $d_{x}$, and $d_{z}$. Each data point displayed is the median across 10 trials of the mean time spent in each portion of the algorithm. When the dependence on $n$ is being considered, $d_{x}$ and $d_{z}$ are both kept at 1. When either $d_{x}$ or $d_{z}$ are being varied, the other is kept at 1 and $n$ is kept at 2000. The first two rows use kernel based functions for both $F$ and $G$ while the third row uses only linear and quadratic functions of $y$ for $G$, which has the same complexity but with lower constants and thus is much faster. These experiments confirm that the complexity of all stages grow linearly with the number of samples and that the descent steps contribute most heavily to the total run time. They also show that, as predicted, the dimension of $x$ increases linearly the time complexity of the main descent steps and the calculations per stage, and the dimension of $z$ increases linearly the complexity of the pre-calculations.
@@ 7 The optimal transport problem
The methodology developed for the OTBP can be easily adapted to solve regular [Monge] OTP. The OTP is simpler than the OTBP, as it involves only two distributions, a source $\rho_{0}$ and a target $\rho_{1}$:
@@
For the purpose of relating them to the barycenter problem, we introduce a binary covariate $z\in\{0,1\}$, and think of the source and target distributions as instances of the single conditional distribution $\rho(x|z)$: $\rho_{0}(x)=\rho(x|z=0),\quad\rho_{1}(x)=\rho(x|z=1)$. Then the map $y=T(x,z)$ that solves the barycenter problem for $\rho(x|z)$, automatically provides the solution to (16) through $Q(x)=T^{-1}\left(T(x,0),1\right)$, a standard result in interpolation displacement [5].
@@
The fact that there is only one, binary covariate $z$ simplifies our methodology considerably, since except for an arbitrary sign, there is only one function $f(z)$ with zero mean and norm one: $f(z)\propto\begin{cases}\ \frac{1}{n_{0}}&\hbox{for }z=0\cr-\frac{1}{n_{1}}&\hbox{for }z=1\end{cases}$. Then the barycenter problem reduces to
@@
Other than the simplifying facts that we do not need to update $a$ and $b$ and the matrix $Q_{z}$ consists of a single column, the procedure to solve (17) follows the same steps as the one for the full barycenter problem (10).
We can bypass the barycenter $\mu$ in the procedure above, finding a map $Q(x)=T(x,0)$ that pushes forward $\rho_{0}$ to $\rho_{1}$ directly, by minimizing $L$ in (17) only over the $y_{i}$ with corresponding $z_{i}=0$, i.e. over $T(x,0)$, leaving the remaining $y_{i}$ fixed at $x$, i.e. setting $T(x,1)=x$. This enforces the condition that $T(x,0)\#\rho_{0}=\rho_{1}$, since the final $y_{i}^{*}=\begin{cases}T\left(x_{i},0\right)&\hbox{for }z_{i}=0\cr x_{i}&\hbox{for }z_{i}=1\end{cases}$ must be independent of $z$. This procedure, while lacking the symmetry of the prior one with respect to $\rho_{0,1}$, is more straightforward, and is particularly well-suited for density estimation.